OpenStudy (squirrels):

Write a recursive definition for the sequence 11, 8, 5, 2... a1=11, an=a(n-1)-4 a1=11, an=a(n-1)-3 a1=2, an=a(n-1)-3 a1=2, an=a(n-1)+3

OpenStudy (squirrels):

\[a _{1}=11, a _{n}=a _{n-1}-4\] Thats what what the options look like in the question above^

OpenStudy (bibby):

Do you see a common subtraction/addition happening from term to term?

OpenStudy (squirrels):

Yes, 3

OpenStudy (bibby):

and what do you think a1 (the first term) is?

OpenStudy (squirrels):

Im guessing 11, but im not too sure

OpenStudy (squirrels):

Or is it -2?

OpenStudy (bibby):

\[\large {5,6,7,8,9} \]\[\large {a_1 = 5} \]\large {a_2 = 6} \] etc.

OpenStudy (bibby):

it's the first term in teh sequence, so 11

OpenStudy (squirrels):

ah i see, i thought so.

OpenStudy (bibby):

so yeah, B. ♥♥

OpenStudy (squirrels):

which leaves us with the answers, a or b?

OpenStudy (squirrels):

OHH!!! that was easy lol thank you bibby ♥♥♥♥♥

OpenStudy (bibby):

it's a1=11, an=a(n-1)-3 as a1 = 11 and every term is the term before it minus 3

OpenStudy (squirrels):

That was too easy bib, i look like such a fool with my ignorance. Would you mind helping me with another one? Last one.

OpenStudy (bibby):

maybe. would you mind "helping me"?

OpenStudy (squirrels):

lmao, it depends

OpenStudy (bibby):

give me a hand buddy boy

OpenStudy (squirrels):

Those were the old days bib, I changed my ways. I'm a follower of Christ now.

OpenStudy (bibby):

No handjobs necessary, post the question bby

OpenStudy (squirrels):

lmao will do.

OpenStudy (squirrels):

Suppose you drop a tennis ball from a height of 9 feet. After the ball hits the floor, it rebounds to 80% of its previous height. How high will the ball rebound after its third bounce? Round to the nearest tenth. 5.8 feet 1 feet 4.6 feet 7.2 feet

OpenStudy (bibby):

The first bounce is 9 The second bounce is 80% of 9. What is 80% of 9?

OpenStudy (squirrels):

7.2..... jesus christ that was easy

OpenStudy (bibby):

not yet

OpenStudy (bibby):

learn to read ytou spic

OpenStudy (bibby):

the third bounce is 80% of 7.2

OpenStudy (squirrels):

lmao im sorry

OpenStudy (squirrels):

and correct

OpenStudy (bibby):

<3. 0.8*7.2=?

OpenStudy (squirrels):

5.76 ♥

OpenStudy (bibby):

which rounds to 5.8 ♥ btw that's a geometric series with the common ratio 80%/0.8 9,7.2,5.76,... etc.

OpenStudy (squirrels):

Ohhh i see, still easy. Sorry homeschooling is a bit challenging for me. I need to be taught by a teacher to learn. I'm not good with self instructions. Thank you bib ♥

OpenStudy (bibby):

I love you dude. Any time ♥

OpenStudy (squirrels):

I'm considering giving you "a hand" now ☺☻♥